S(with a sideways 8 beside it)=8-2+1/2+...=? Please help! I am so horrible at this!
Answer by fractalier(6550)

Yes, okay...this is an infinite sum with the first term equalling 8 and the common ratio being -1/4...
The formula for this is
a-sub-1 / (1 - r)
So plugging in we get...
8 / (1 - (-1/4)) =
8 / (5/4) =
32/5